Obtain the general solutions given the differential operators (D + 3)4 y = 0

1
Expert's answer
2022-01-17T18:07:12-0500

Solution:

"4(D + 3) y = 0
\\ \Rightarrow (D + 3) y = 0
\\ \Rightarrow (\dfrac d{dx} + 3) y = 0
\\ \Rightarrow \dfrac {dy}{dx}+3y=0
\\ \Rightarrow \dfrac {dy}{dx}=-3y
\\ \Rightarrow \dfrac {dy}{y}=-3dx"

On integrating both sides,

"\log y=-3x+C
\\ \Rightarrow y=e^{-3x+C}
\\ \Rightarrow y=C_1e^{-3x}, \ where \ C_1=e^C"